OpenStudy (anonymous):

got one more 4 1/5divided by 3 1/7

OpenStudy (anonymous):

4 1/5 = 21/5 3 1/7 = 22/7 so (21/5)/(22/7) which is (21*7)/(5*22) = 147/110 = 1 37/110